# Efficient Subdivision of Bézier Curves/Surfaces via Blossoms

### v1

@@ 2.1 Subdivision of a Bézier Curve
For any polynomial curve $C(u)=\sum_{i=0}^{n}{\bf c}_{i}u^{i}$ of degree $n$, $n\in\mathbb{N}$, ${\bf c}_{i}\in\mathbb{R}^{3}$, $i=0,\dots,n$, the control points ${\bf w}_{\nu}$, $\nu=0,\dots,n$, of the Bézier curve $C^{1}$ on $C$ defined for $a\leq u\leq b$, $a,b\in\mathbb{R}$, are (see (Goldman, 2003))
${\bf w}_{\nu}=b\tilde{\ }(\underbrace{b,\dots,b}_{\text{$\nu$}},\underbrace{a,\dots,a}_{\text{$n-\nu$}}),$ (1)
where $b\tilde{\ }(u_{1},\dots,u_{n})=\sum_{i=0}^{n}{\bf c}_{i}b_{i}\tilde{\ }(u_{1},\dots,u_{n})$ is the blossom of $C(u)$, and
@@
is the blossom of the monomial $u^{i}$. The sums in (2) are taken over all distinct indices $\{\alpha_{1},\dots\alpha_{i}\}$. In the case where $i=0$ we have $b_{0}\tilde{\ }(u_{1},\dots,u_{n})=1$. For example, if $C(u)$ is of degree 2 then $b_{0}\tilde{\ }(u_{1},u_{2})=1$, $b_{1}\tilde{\ }(u_{1},u_{2})=\frac{u_{1}+u_{2}}{2}$ and $b_{2}\tilde{\ }(u_{1},u_{2})=u_{1}u_{2}$. We note that formula (2) in this form is not easy to be implemented. In the next theorem we establish formula (1) in an equivalent closed form that is more suitable for computations.
@@
Proof. According to (2), we have
@@ 2.2 Subdivision of a Tensor Product Bézier Surface
For any polynomial surface $S(u,v)=\sum_{i=0}^{n}\sum_{j=0}^{m}{\bf c}_{ij}u^{i}v^{j}$ of degree $(n,m)$, $n,m\in\mathbb{N}$, ${\bf c}_{ij}\in\mathbb{R}^{3}$, $i=0,\dots,n$, $j=0,\dots,m$, the control points ${\bf p}_{\nu\mu}$, $\nu=0,\dots,n$, $\mu=0,\dots,m$, of the TPB surface patch $S^{1}$ on $S$ defined for $a\leq u\leq b$ and $c\leq v\leq d$, are (see (Goldman, 2003), pp. 338-339)
${\bf p}_{\nu\mu}=b^{\Box}(\underbrace{b,\dots,b}_{\text{$\nu$}},\underbrace{a,\dots,a}_{\text{$n-\nu$}},\underbrace{d,\dots,d}_{\text{$\mu$}},\underbrace{c,\dots,c}_{\text{$m-\mu$}}),$ (6)
where
$b^{\Box}(u_{1},\dots,u_{n},v_{1},\dots,v_{m})=\sum_{i=0}^{n}\sum_{j=0}^{m}{\bf c}_{ij}b_{ij}^{\Box}(u_{1},\dots,u_{n},v_{1},\dots,v_{m})$
is the blossom of $S(u,v)$, and
@@
is the blossom of the monomial $u^{i}v^{j}$. The sums in (7) are taken over all distinct indices $\{\alpha_{1},\dots\alpha_{i}\}$ and all distinct indices $\{\beta_{1},\dots\beta_{i}\}$. In the case where either $i=0$, or $j=0$, the corresponding sum in (7) equals 1. For example, if $S(u,v)$ is of degree $(3,3)$ then the blossom of $u^{2}v$ is $b_{21}^{\Box}(u_{1},u_{2},u_{3},v_{1},v_{2},v_{3})=\frac{u_{1}u_{2}+u_{2}u_{3}+u_{1}u_{3}}{3}\cdot\frac{v_{1}+v_{2}+v_{3}}{3}$ and the blossom of $v^{3}$ is $b_{03}^{\Box}(u_{1},u_{2},u_{3},v_{1},v_{2},v_{3})=v_{1}v_{2}v_{3}$. Again, we note that formula (7) in this form is not easy to be implemented. Therefore, next we establish formula (6) in an equivalent closed form that is more suitable for computations.
@@
Proof. According to (7), we have
$b_{ij}^{\Box}(u_{1},\dots,u_{n},v_{1},\dots,v_{m}):=\frac{1}{\binom{n}{i}\binom{m}{j}}\sum_{\bar{\alpha}_{i}\in{\cal A}_{i}}U(\bar{\alpha}_{i})\sum_{\bar{\beta}_{j}\in{\cal B}_{j}}V(\bar{\beta}_{j})$
where ${\cal A}_{i}$ is the set of all subsets $\bar{\alpha}_{i}:=\{\alpha_{1},\dots,\alpha_{i}\}$ of distinct indices of size $i$, ${\cal B}_{j}$ is the set of all subsets $\bar{\beta}_{j}:=\{\beta_{1},\dots,\beta_{j}\}$ of distinct indices of size $j$, $U(\bar{\alpha}_{i}):=u_{\alpha_{1}}\dots u_{\alpha_{i}}$, and $V(\bar{\beta}_{j}):=v_{\beta_{1}}\dots v_{\beta_{j}}$.
Next, we will compute $b_{ij}^{\Box}(\underbrace{b,\dots,b}_{\text{$\nu$}},\underbrace{a,\dots,a}_{\text{$n-\nu$}},\underbrace{d,\dots,d}_{\text{$\mu$}},\underbrace{c,\dots,c}_{\text{$m-\mu$}})$. For a fixed $\bar{\alpha}_{i}$ from Theorem 1 we have
Analogously, we obtain
$\sum_{\bar{\beta}_{j}\in{\cal B}_{j}}V(\bar{\beta}_{j})=\sum_{r=\max(0,j+\mu-m)}^{\min(j,\mu)}\binom{\mu}{r}\binom{m-\mu}{j-r}d^{r}c^{j-r}.$ (10)
By substitution of (9) and (10) in (7) and (6) we get
@@ 2.3 Subdivision of a Triangular Bézier Surface
Let ${\bf a}=(a_{1},a_{2})$, ${\bf b}=(b_{1},b_{2})$, ${\bf c}=(c_{1},c_{2})$ be three non-collinear points in the plane $Oxy$. For any polynomial surface $S(u,v)=\sum_{i=0}^{n}\sum_{j=0}^{m}{\bf c}_{ij}u^{i}v^{j}$ of total degree $N:=n+m$, $n,m\in\mathbb{N}$, ${\bf c}_{ij}\in\mathbb{R}^{3}$, $i=0,\dots,n$, $j=0,\dots,m$, the control points ${\bf q}_{\nu\mu}$, $\nu=0,\dots,N$, $\mu=0,\dots,N-\nu$, of the TB surface patch $S^{2}$ defined in the triangle $\triangle\bf{abc}$, are (see (Goldman, 2003), pp. 338-339)
${\bf q}_{\nu\mu}=b^{\triangle}(\underbrace{{\bf a},\dots,{\bf a}}_{\nu},\underbrace{{\bf b},\dots,{\bf b}}_{\mu},\underbrace{{\bf c},\dots\bf c}_{N-\nu-\mu}),$ (11)
where
$b^{\triangle}((u_{1},v_{1}),\dots,(u_{N},v_{N}))=\sum_{i=0}^{n}\sum_{j=0}^{m}{\bf c}_{ij}b_{ij}^{\triangle}((u_{1},v_{1}),\dots,(u_{N},v_{N}))$
is the blossom of $S(u,v)$, and
@@
$\binom{N}{i,j}=\frac{N!}{i!j!(N-i-j)!}$, is the blossom of the monomial $u^{i}v^{j}$. The sum in (12) is taken over all sets of distinct indices $\{\alpha_{1},\dots,\alpha_{i},\beta_{1},\dots,\beta_{j}\}$. In the case where either $i=0$, or $j=0$, the corresponding sum in (12) becomes $\sum_{\scriptstyle\{\alpha_{1},\dots,\alpha_{i}\}\subseteq\{1,\dots,N\}}\frac{u_{\alpha_{1}}\dots u_{\alpha_{i}}}{\binom{N}{j}}$ and $\sum_{\scriptstyle\{\beta_{1},\dots,\beta_{j}\}\subseteq\{1,\dots,N\}}\frac{v_{\beta_{1}}\dots v_{\beta_{j}}}{\binom{N}{i}}$, respectively. For example, if $S(u,v)$ is of degree $(3,3)$ then the blossom of $u^{2}v$ is $b_{21}^{\triangle}(u_{1},u_{2},u_{3},v_{1},v_{2},v_{3})=\frac{u_{1}u_{2}v_{3}+u_{2}u_{3}v_{1}+u_{1}u_{3}v_{2}}{3}$ and the blossom of $v^{3}$ is $b_{03}^{\triangle}(u_{1},u_{2},u_{3},v_{1},v_{2},v_{3})=v_{1}v_{2}v_{3}$. Again, we note that formula (12) in this form is not easy to be implemented. Therefore, similar to the case of TPB surfaces, in the next theorem we present formula (11) in an equivalent closed form that is more suitable for evaluations of the control points.

### v2

@@ 2.1 Subdivision of a Bézier Curve
Let $C(u)=\sum_{i=0}^{n}{\bf c}_{i}u^{i}$ be a polynomial curve of degree $n$, $n\in\mathbb{N}$, ${\bf c}_{i}\in\mathbb{R}^{3}$, $i=0,\dots,n$. Let the Bézier curve $C^{1}(u)=\sum_{\nu=0}^{n}{\bf w}_{\nu}B_{\nu}^{n}(u)$ be the part of $C(u)$ defined in the interval $[a,b]$, where $a,b\in\mathbb{R}$, ${\bf w}_{\nu}$ are the control points of $C^{1}$, and $B_{\nu}^{n}(u)=\binom{n}{\nu}u^{\nu}(1-u)^{n-\nu}$ are the Bernstein polynomials, $\nu=0,\dots,n$.
Let $b\tilde{\ }(u_{1},\dots,u_{n})$ be the blossom of $C(u)$. It is well known that
${\bf w}_{\nu}=b\tilde{\ }(\underbrace{b,\dots,b}_{\text{$\nu$}},\underbrace{a,\dots,a}_{\text{$n-\nu$}}).$ (1)
Furthermore, for the blossom $b\tilde{\ }$ we have the formula
$b\tilde{\ }(u_{1},\dots,u_{n})=\sum_{i=0}^{n}{\bf c}_{i}b_{i}\tilde{\ }(u_{1},\dots,u_{n}),\ {\rm where}$
@@
is the blossom of the monomial $u^{i}$. The sums in (2) are taken over all distinct indices $\{\alpha_{1},\dots\alpha_{i}\}$ and $b_{0}\tilde{\ }(u_{1},\dots,u_{n})=1$.
We have observed that the evaluation of blossoms using formula (2) is not straightforward since it requires the generation of all subsets $\{\alpha_{1},\dots,\alpha_{i}\}$. This make the evaluation of control points computationally intense. In the next theorem we propose a formula for direct computation of the control points ${\bf w}_{\nu}$, $\nu=0,\dots,n$.
@@
Proof According to (2), we have
@@ 2.2 Subdivision of a Tensor Product Bézier Surface
Let $S(u,v)=\sum_{i=0}^{n}\sum_{j=0}^{m}{\bf c}_{ij}u^{i}v^{j}$ be a polynomial surface of bidegree $(n,m)$, $n,m\in\mathbb{N}$, ${\bf c}_{ij}\in\mathbb{R}^{3}$, $i=0,\dots,n$, $j=0,\dots,m$. Let the TPB surface $S^{1}(u,v)=\sum_{\nu=0}^{n}\sum_{\mu=0}^{m}{\bf p}_{\nu\mu}B_{\nu}^{n}(u)B_{\mu}^{n}(v)$ be the part of $S(u,v)$ defined for $a\leq u\leq b$ and $c\leq v\leq d$, where $a,b,c,d\in\mathbb{R}$ and ${\bf p}_{\nu\mu}$ are the control points of $S^{1}$, $\nu=0,\dots,n$, $\mu=0,\dots,m$.
Let $b^{\Box}(u_{1},\dots,u_{n},v_{1},\dots,v_{m})$ be the blossom of $S(u,v)$. It is known that
${\bf p}_{\nu\mu}=b^{\Box}(\underbrace{b,\dots,b}_{\text{$\nu$}},\underbrace{a,\dots,a}_{\text{$n-\nu$}},\underbrace{d,\dots,d}_{\text{$\mu$}},\underbrace{c,\dots,c}_{\text{$m-\mu$}}).$ (6)
Furthermore, for the blossom $b^{\Box}$ we have the formula (see (Goldman, 2003), pp. 338-339)
$b^{\Box}(u_{1},\dots,u_{n},v_{1},\dots,v_{m})=\sum_{i=0}^{n}\sum_{j=0}^{m}{\bf c}_{ij}b_{ij}^{\Box}(u_{1},\dots,u_{n},v_{1},\dots,v_{m}),\ \rm{where}$ (7)
@@
is the blossom of the monomial $u^{i}v^{j}$. The sums in (8) are taken over all distinct indices $\{\alpha_{1},\dots\alpha_{i}\}$ and all distinct indices $\{\beta_{1},\dots\beta_{i}\}$. In the case where either $i=0$, or $j=0$, the corresponding sum in (8) equals 1.
We note that formula (8) in this form is not easy to be implemented. Therefore, next we establish formula (6) in an equivalent closed form that is more suitable for computations.
@@
Proof We observe that formula (8) can be presented in the equivalent form
$b_{ij}^{\Box}(u_{1},\dots,u_{n},v_{1},\dots,v_{m})=b_{i}\tilde{\ }(u_{1},\dots,u_{n})b_{j}\tilde{\ }(v_{1},\dots,v_{m}),$ (10)
where $b_{i}\tilde{\ }$ and $b_{j}\tilde{\ }$ are the blossoms of the monomials $u^{i}$ and $v^{j}$, respectively. From Theorem 1 we have
$b_{i}\tilde{\ }(u_{1},\dots,u_{n})=\frac{1}{\binom{n}{i}}\sum_{k=\max(0,i+\nu-n)}^{\min(i,\nu)}\binom{\nu}{k}\binom{n-\nu}{i-k}b^{k}a^{n-k},$
$b_{j}\tilde{\ }(v_{1},\dots,v_{m})=\frac{1}{\binom{m}{j}}\sum_{r=\max(0,j+\mu-m)}^{\min(j,\mu)}\binom{\mu}{r}\binom{m-\mu}{j-r}d^{r}c^{j-r}.$
Substituting (11) in (10) and (7), we obtain from (6)
@@ 2.3 Subdivision of a Triangular Bézier Surface
Let $S(u,v)=\sum_{i=0}^{n}\sum_{j=0}^{m}{\bf c}_{ij}u^{i}v^{j}$ be a polynomial surface of total degree $N=n+m$, $n,m\in\mathbb{N}$, ${\bf c}_{ij}\in\mathbb{R}^{3}$, $i=0,\dots,n$, $j=0,\dots,m$, and ${\bf a}=(a_{1},a_{2})$, ${\bf b}=(b_{1},b_{2})$, ${\bf c}=(c_{1},c_{2})$ be three non-collinear points in $\mathbb{R}^{2}$. Let the TB surface
$S^{2}(u,v)=\sum_{\nu=0}^{N}\sum_{\mu=0}^{N-\nu}{\bf q}_{\nu\mu}B_{\nu\mu}^{N}(u,v)u^{\nu}v^{\mu}(1-u-v)^{N-\nu-\mu}$
be the part of $S(u,v)$ defined in the triangle $\triangle\bf{abc}$, where ${\bf q}_{\nu\mu}$ are the control points of $S^{2}$,
$B_{\nu\mu}^{N}(u,v)=\binom{N}{\nu,\mu}u^{\nu}v^{\mu}(1-u-v)^{N-\nu-\mu},\ \nu=0,\dots,N,\ \mu=0,\dots,N-\nu$
are the bivariate Bernstein polynomials, and $\binom{N}{\nu,\mu}=\frac{N!}{\nu!\mu!(N-\nu-\mu)!}$.
Let $b^{\triangle}((u_{1},v_{1}),\dots,(u_{N},v_{N}))$ be the blossom of $S(u,v)$. It is known that
${\bf q}_{\nu\mu}=b^{\triangle}(\underbrace{{\bf a},\dots,{\bf a}}_{\nu},\underbrace{{\bf b},\dots,{\bf b}}_{\mu},\underbrace{{\bf c},\dots\bf c}_{N-\nu-\mu}).$ (12)
Furthermore, for the blossom $b^{\triangle}$ we have the formula (see (Goldman, 2003), pp. 338-339)
$b^{\triangle}((u_{1},v_{1}),\dots,(u_{N},v_{N}))=\sum_{i=0}^{n}\sum_{j=0}^{m}{\bf c}_{ij}b_{ij}^{\triangle}((u_{1},v_{1}),\dots,(u_{N},v_{N})),\ \rm{where}$
@@
is the blossom of the monomial $u^{i}v^{j}$. The sum in (13) is taken over all sets of distinct indices $\{\alpha_{1},\dots,\alpha_{i},\beta_{1},\dots,\beta_{j}\}$ and $b_{i0}^{\triangle}((u_{1},v_{1}),\dots,(u_{N},v_{N}))=b_{i}\tilde{\ }(u_{1},\dots,u_{N})$, $b_{0j}^{\triangle}((u_{1},v_{1}),\dots,(u_{N},v_{N}))=b_{j}\tilde{\ }(v_{1},\dots,v_{N}).$
We note that the evaluation of blossoms using formula (13) is not straightforward and easy for implementation. Therefore, similar to the case of TPB surfaces, in the next theorem we propose a formula for direct computation of the control points ${\bf q}_{\nu\mu}$, $\nu=0,\dots,N$, $\mu=0,\dots,N-\nu$.